OpenStudy (anonymous):

Find a basis that includes u = {(1,5,4)} in R^3.

OpenStudy (helder_edwin):

any base?

OpenStudy (helder_edwin):

the easy way or the not-so-hard way?

OpenStudy (anonymous):

that's the only given. both ways will do :)

OpenStudy (helder_edwin):

do u know about inner products?

OpenStudy (helder_edwin):

if so then just solve \[ \large 0=(x,y,z)\cdot(1,5,4)=x+5y+4z \]

OpenStudy (anonymous):

we haven't discussed about the inner products or maybe I just don't know anything about it. Its our quiz tomorrow on vector spaces, spanning, linear combination,linear dependence and independence, and basis. Got any tips? :)

OpenStudy (sirm3d):

choose a 2nd that is most convenient, say (1,0,0), then work on the third basis (a,b,c) using the first two (1,5,4) and (1,0,0).

OpenStudy (helder_edwin):

yes. just pick any vector not parallel (a scalar multiple) to the one given. any vector

OpenStudy (helder_edwin):

then u have two choices: (i) use the cross product u learned in multi-variable calculus to get the third vector, for instance if u choose (1,0,0) then \[ \large (1,5,4)\times(1,0,0)= \] is the third one

OpenStudy (helder_edwin):

(ii) check the consistency condition for the system \[ \large \left(\begin{array}{cc|c} 1 & 1 & a\\ 5 & 0 & b\\ 4 & 0 & c \end{array}\right) \]

OpenStudy (helder_edwin):

this last one is more suited with what your examination is going to be about.

OpenStudy (helder_edwin):

then pick a vector that doesn't fullfil the condition